英語 (English)

基盤英語 (Basic English) (工 ((機 B)1 年))

ラックストンロバート · 非常勤講師 1 単位前期水 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】This class will stress English basics, including grammar and pronunciation, as well as give students a chance to improve their conversational ability

【授業の概要】We will do some TOEIC practice, textbook exercises, role-plays

and skits. 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students will increase their confidence about speaking English. 【授業の計画】

1. Introduction

2. Unit 1.1

3. Unit 1.2

4. Unit 2.1

5. Unit 2.2

6. Unit 3.1

7. Quiz 1 and Skit Preparation

8. Skit Day

9. Unit 3.2

10. .Unit 4.1

11. Unit 4.2

12. Unit 5.1

13. Unit 5.2

14. TOEIC Practice Day

15. Quiz 2 and Skit Preparation

16. Skit Day 【教科書】 Jack C. Richards: New Person to Person, Student Book 1 【参考書等】 [参考資料] 【成績評価の方法】 Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% 【再試験の有無】 [再評価] 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220847 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ラックストン .

English

Basic English (工 ((機 B)1 年))

Robert Luxton · Part-time Lecturer 1 unit 前期水 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ This class will stress English basics, including grammar and pronunciation,

as well as give students a chance to improve their conversational ability Outline⟩ We will do some TOEIC practice, textbook exercises, role-plays and

skits. �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students will increase their confidence about speaking English.

Schedule⟩

1. Introduction

2. Unit 1.1

3. Unit 1.2

4. Unit 2.1

5. Unit 2.2

6. Unit 3.1

7. Quiz 1 and Skit Preparation

8. Skit Day

9. Unit 3.2

10. .Unit 4.1

11. Unit 4.2

12. Unit 5.1

13. Unit 5.2

14. TOEIC Practice Day

15. Quiz 2 and Skit Preparation

16. Skit Day Textbook⟩ Jack C. Richards: New Person to Person, Student Book 1 Reference⟩ [参考資料] Evaluation Criteria⟩ Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% Re-evaluation⟩ [再評価] Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220847 Contact (Office-Hour, Room, E-mail)⟩

⇒ Luxton .